# DragText: Rethinking Text Embedding in Point-based Image Editing

### v1

@@ 4.2.1 Motion Supervision
Motion supervision drags image content around handle points $h_{i}^{k}$ toward target points $g_{i}$ on the image feature space by optimizing the latent vector $\mathbf{\hat{z}}^{k}_{t}$, where $k$ implies the number of drag editing iterations. The feature map $\mathcal{F}(\mathbf{\hat{z}}^{k}_{t},\mathbf{\hat{c}}^{k})$ is obtained from the third decoder block of the denoising U-Net. Its corresponding feature vector at the specific pixel location $q$ is denoted as $\mathcal{F}_{q}(\mathbf{\hat{z}}^{k}_{t},\mathbf{\hat{c}}^{k})$.
@@
$\small\mathcal{L}_{\text{ms}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+d_{i}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k}))\right\|_{1}$ (1)
$+\lambda_{\text{image}}\left\|\left(\mathbf{\hat{z}}_{t-1}^{k}-\text{sg}(\mathbf{\hat{z}}_{t-1}^{0})\right)\odot(\mathbbm{1}-M_{\text{image}})\right\|_{1},$
where $\text{sg}(\cdot)$ is the stop-gradient operator. It prevents the gradient from being backpropagated for the term $\text{sg}(\cdot)$. $q_{1}=\Omega(h_{i}^{k},r_{1})=\{(x,y):|x-x_{i}^{k}|\leq r_{1},|y-y_{i}^{k}|\leq r_{1}\}$ describes the square region centered at $h_{i}^{k}$ with radius $r_{1}$. $d_{i}=\left(g_{i}-h_{i}^{k}\right)/\left\|g_{i}-h_{i}^{k}\right\|_{2}$ is the normalized vector from $h^{k}_{i}$ to $g_{i}$, thus $\mathcal{F}_{q_{1}+d_{i}}$ is computed through the bilinear interpolation. This design allows $h^{k}_{i}$ to be moved to $h^{k}_{i}+d_{i}$, but not the reverse with $\text{sg}(\cdot)$. $M_{\text{image}}$ is the binary mask drawn by the user to define an editable region. The second term of the equation adjusts the extent to which regions outside the mask remain unchanged during the latent optimization process with $\lambda_{\text{image}}$.
For each iteration $k$, $\mathbf{\hat{z}}_{t}^{k}$ undergoes a gradient descent step to minimize $\mathcal{L}_{\text{ms}}$:
$\small\mathbf{\hat{z}}_{t}^{k+1}=\mathbf{\hat{z}}_{t}^{k}-\eta_{\text{ms}}\cdot\frac{\partial\mathcal{L}_{\text{ms}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})}{\partial\mathbf{\hat{z}}_{t}^{k}},$ (2)
@@ 4.2.2 Text Optimization
The input text is encoded to the text embedding $\mathbf{c}\in\mathbb{R}^{77\times d}$ through the CLIP text encoder [22]. $77$ represents the number of text tokens, out of which $l$ tokens carry semantic meaning. Here, the $l$ tokens represent the number of tokens obtained when the text passes through the tokenizer, typically calculated as the number of words plus one. To optimize the text embedding $\mathbf{\hat{c}}^{k}$, we use the same loss function as in the motion supervision. It allows $\mathbf{\hat{c}}^{k}$ to follow the latent optimization towards the dragging direction. In addition, to preserve important content in the original text embedding $\mathbf{c}^{0}$, we use the mask $M_{\text{text}}\in\mathbb{R}^{l\times d}$ and regularize the text optimization process.
@@
$\small\mathcal{L}_{\text{text}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+d_{i}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k}))\right\|_{1}$ (3)
$+\lambda_{\text{text}}\left\|\left(\mathbf{\hat{c}}^{k}-\text{sg}(\mathbf{c}^{0}))\right)\odot M_{\text{text}}\right\|_{1}$
For each iteration, $\mathbf{\hat{c}}^{k}$ undergoes a gradient descent step to minimize $L_{\text{text}}$:
$\small\mathbf{\hat{c}}^{k+1}=\mathbf{\hat{c}}^{k}-\eta_{\text{text}}\cdot\frac{\partial\mathcal{L}_{\text{text}}(\mathbf{\hat{z}}_{t}^{k},\mathbf{\hat{c}}^{k})}{\partial\mathbf{\hat{c}}^{k}},$ (4)
@@ 4.2.3 Point Tracking
After the latent vector $\mathbf{\hat{z}}_{t}^{k}$ and text embedding $\mathbf{\hat{c}}^{k}$ are optimized at the $k$-th iteration step, the positions of the handle point $h_{i}^{k}$ change according to the content dragging. Thus, it is necessary to track the new handle point $h_{i}^{k+1}$ in the updated feature map $\mathcal{F}(\mathbf{\hat{z}}_{t}^{k+1},\mathbf{\hat{c}}^{k+1})$. To find the new handle points $h_{i}^{k+1}$ corresponding to the previous handle points $h_{i}^{k}$, we find the region in the feature map $\mathcal{F}(\mathbf{\hat{z}}_{t}^{k+1},\mathbf{\hat{c}}^{k+1})$ that is most similar to the region around the initial handle points $h_{i}^{0}$ in the feature map $\mathcal{F}(\mathbf{z}_{t}^{0},\mathbf{\hat{c}}^{0})$:
$\small h_{i}^{k+1}=\mathop{\arg\min}_{q_{2}}\left\|F_{q_{2}}(\mathbf{\hat{z}}_{t}^{k+1},\mathbf{\hat{c}}^{k+1})-F_{h_{i}^{0}}(\mathbf{z}^{0}_{t},\mathbf{c}^{0})\right\|_{1}$ (5)
where $q_{2}$ denotes a point within the region $\Omega(h_{i}^{k},r_{2})$. This ensures that handle points are consistently updated via nearest neighbor search for subsequent optimization iterations.
@@ 5 Experiments
To evaluate the effectiveness of the proposed DragText, we conduct a series of experiments on DragDiffusion [29], FreeDrag [12], DragNoise [13], and GoodDrag [40]. DragText is applicable to these methods without hyperparameter tuning, ensuring robustness. Moreover, hyperparameters for text optimization are uniformly applied across all methods.
Specifically, we set $\lambda_{\text{text}}=0.1$ to match the image regularization factor $\lambda_{\text{image}}=0.1$, and the text optimization learning rate $\eta_{\text{text}}=0.004$. Our method operates without utilizing additional GPU resources beyond baseline methods and is executed using a single NVIDIA RTX A6000 GPU. Qualitative results are based on DragDiffusion, and further examples using other models are in the Supplement.
5.1 Qualitative Evaluation
Table 1: Quantitative results on the DragBench dataset
@@ 5.1 Qualitative Evaluation
Figure 7: Quantitative results comparing the product of LPIPS and MD for baseline models and w/ DragText. Quantitative performance improves across all models with DragText.
Figure 8: Ablation study on the number of U-Net decoder blocks, determining the location where the feature map is obtained. Lower U-Net blocks (e.g., Block 1) focus on low semantic information, while higher U-Net blocks (e.g., Block 4) concentrate on high semantic information. The best results are observed in block 3, where both types of information are appropriately mixed.
In Fig. 5, we compared qualitative results with and without DragText. Existing methods struggle to reach target points or suffer from the degradation of important image semantics, due to the use of image and text embeddings that are not strongly coupled. In contrast, by applying DragText, images can be accurately dragged to the specified target points (e.g., (b) mountain peaks and (d) statues) while preserving essential semantics that should not disappear (e.g., (c) the mouth of a crocodile). Additionally, DragText effectively manages content removal and creation caused by dragging, producing high-quality results.
@@ 5.2 Quantitative Evaluation
In Table 1, we present the quantitative results on the DragBench dataset [29] for the four baseline methods and methods with DragText applied to these baselines. The evaluation metrics are LPIPS [38] and mean distance (MD) [21]. LPIPS measures the perceptual similarity to the original image as a perceptual loss. MD measures the distance between the final position of the handle point $h$ and the target point $g$ in the feature map of the DIFT model [33]. There is little difference in performance between the baseline models and the DragText applied models in terms of LPIPS, but there is a significant improvement in MD. As shown in Fig. 7, this difference is more dramatic. It shows that applying DragText results in substantial performance improvements across all models. This indicates that text embedding optimization allows for effective dragging while maintaining image fidelity.
@@ 5.3 Ablation Study
Effect of Text Regularization. In Fig. 6, we investigated the impact of $\lambda_{\text{text}}$ designed to preserve the semantics of the image. We edited images by varying $\lambda_{\text{text}}$ from 0 to 10 to regularize the text embedding. As $\lambda_{\text{text}}$ approaches 0, the force of dragging becomes more dominant than the preservation of semantics. For example, in Fig. 6 (d), the icecap has disappeared, indicating extensive dragging. Conversely, applying an excessively high $\lambda_{\text{text}}$ maintains the original content of the image and inhibits editing during text embedding optimization. In Fig. 6 (a), minimal dragging is observed, showing that the original image remains largely unchanged. We found that setting $\lambda_{\text{text}}$ to 0.1 achieves the optimal balance between dragging and content preservation, and it can be changed by users.
Figure 9: Changes in images when manipulating the image and text embeddings. $\omega$ is a factor of the linear interpolation between the original embeddings and the dragged embeddings. As $\omega$ approaches 0, the image resembles the original image; as $\omega$ approaches 1, the image resembles the dragged image. DragText can also generate images outside the originally intended editing range (e.g., extrapolating to $\omega=-1.0$ and $\omega=2)$ in the direction of the drag preserving the semantic content at the same time.
Effect of the Number of U-Net Decoder Block. In Fig. 8, we perform DragText using feature maps from four different U-Net decoder blocks. To assess the impact of U-Net decoder blocks on text embedding, the image embedding is fixed using the feature map from the 3rd block. The images optimized with the feature map from the 3rd block exhibited optimal semantic preservation and effective dragging. Feature maps from lower blocks (e.g., Block 1) make it difficult to maintain semantics, whereas feature maps from higher blocks (e.g., Block 4) preserve semantics well but result in poor dragging performance. This phenomenon is likely due to the lower blocks of the U-Net containing low-frequency information of the image [30]. Quantitative metrics supporting our qualitative evaluation are in the Supplement.
@@ 6 Manipulating Embeddings
DragText enables the model to comprehend the transformation and degree of changes in the image while preserving the important semantic content. In Fig. 9, we applied the same linear interpolation to both the image and text embeddings before and after editing to generate the intermediate images. The image and text embeddings for intermediate images are defined as follows: $\mathbf{z}=(1-\omega)\mathbf{z_{0}}+\omega\mathbf{\hat{z}}_{t}^{k},\quad\mathbf{c}=(1-\omega)\mathbf{c^{0}}+\omega\mathbf{\hat{c}}^{k}.$ When $\omega=0$, the result corresponds to the original image, and as $\omega$ approaches 1, it gets closer to the dragged image. The remarkable point is that this phenomenon is maintained even for $\omega\notin[0,1]$. This allows for adjusting the degree of dragging even beyond the editing process. By applying DragText, even with simple manipulations, the semantic content is preserved while controlling the degree and direction of the drag. This is possible because DragText optimizes the text embedding and image embedding together. In contrast, the baseline model fails to maintain semantics as the text embedding necessary for semantic preservation is not jointly optimized.
@@ 7 Limitation
Diffusion-based editing models have the common problem of often losing the dragged feature in the latent space, leading to the vanishing of content in surrounding objects that are not the editing target (Fig. 10 (a)). For instance, grape grains within the red circle disappear when images are dragged, regardless of whether DragText is applied. However, this situation changes dramatically when the user provides a more detailed text prompt. For instance, in Fig. 10 (b), although the user provides the exact word grapes, the baseline model still cannot drag the grape feature. On the other hand, DragText is able to drag the grape feature because it optimizes the text embedding along with the image embedding, preserving the original content. The optimized text embedding plays a complementary role for the features lost in the image embedding space.
Figure 10: Limitations of diffusion-based editing models. (a) The diffusion-based model sometimes loses features during the drag optimization process, leading to situations where the content is not preserved. However, as shown in (b), when DragText is applied, the drag is performed well while simultaneously preserving the content, unlike the baseline model, if the user provides content information that needs to be dragged.

### v2

@@ 4.2.1 Motion Supervision
Motion supervision drags image content around handle points $h_{i}^{k}$ toward target points $g_{i}$ on the image feature space by optimizing the latent vector $\hat{\mathbf{z}}^{k}_{t}$, where $k$ implies the number of drag editing iterations. The feature map $\mathcal{F}(\hat{\mathbf{z}}^{k}_{t},\hat{\mathbf{c}}^{k})$ is obtained from the third decoder block of the denoising U-Net, and its corresponding feature vector at the specific pixel location $q$ is denoted as $\mathcal{F}_{q}(\hat{\mathbf{z}}^{k}_{t},\hat{\mathbf{c}}^{k})$.
@@
$\mathcal{L}_{\text{ms}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+d_{i}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k}))\right\|_{1}$ (1)
$+\lambda_{\text{image}}\left\|\left(\hat{\mathbf{z}}_{t-1}^{k}-\text{sg}(\hat{\mathbf{z}}_{t-1}^{0})\right)\odot(\mathbbm{1}-M_{\text{image}})\right\|_{1},$
where $\text{sg}(\cdot)$ is the stop-gradient operator to prevent the gradient from being backpropagated. $q_{1}=\Omega(h_{i}^{k},r_{1})=\{(x,y):|x-x_{i}^{k}|\leq r_{1},|y-y_{i}^{k}|\leq r_{1}\}$ describes the square region centered at $h_{i}^{k}$ with radius $r_{1}$. $d_{i}=\left(g_{i}-h_{i}^{k}\right)/\left\|g_{i}-h_{i}^{k}\right\|_{2}$ is the normalized vector from $h^{k}_{i}$ to $g_{i}$ which enables $\mathcal{F}_{q_{1}+d_{i}}$ to be computed by the bilinear interpolation. The first term allows $h^{k}_{i}$ to be moved to $h^{k}_{i}+d_{i}$, but not the reverse with $\text{sg}(\cdot)$. $M_{\text{image}}$ is the binary mask drawn by the user defining an editable region. The second term adjusts the extent to which regions outside the mask remain unchanged with $\lambda_{\text{image}}$ during the optimization.
For each iteration $k$, $\hat{\mathbf{z}}_{t}^{k}$ undergoes a gradient descent step to minimize $\mathcal{L}_{\text{ms}}$:
$\hat{\mathbf{z}}_{t}^{k+1}=\hat{\mathbf{z}}_{t}^{k}-\eta_{\text{ms}}\frac{\partial\mathcal{L}_{\text{ms}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})}{\partial\hat{\mathbf{z}}_{t}^{k}},$ (2)
@@ 4.2.2 Text Optimization
The input text is encoded to the text embedding $\mathbf{c}\in\mathbb{R}^{77\times d}$ through the CLIP text encoder [24]. $77$ represents the total number of text tokens, obtained when text is processed by the tokenizer, with only $l$ tokens carrying semantic meaning. We use the same loss function as in the motion supervision to optimize the text embedding $\mathbf{\hat{c}}^{k}$ for ensuring consistency and generalizability. It allows $\hat{\mathbf{c}}^{k}$ to follow the latent optimization towards the dragging direction. In addition, to preserve important content in the original text embedding $\mathbf{c}^{0}$, we use the mask $M_{\text{text}}\in\mathbb{R}^{l\times d}$ and regularize the text optimization process.
@@
$\mathcal{L}_{\text{text}}=$ $\sum_{i=1}^{n}\sum_{q_{1}}\left\|\mathcal{F}_{q_{1}+d_{i}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})-\text{sg}(\mathcal{F}_{q_{1}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k}))\right\|_{1}$ (3)
$+\lambda_{\text{text}}\left\|\left(\hat{\mathbf{c}}^{k}-\text{sg}(\mathbf{c}^{0}))\right)\odot M_{\text{text}}\right\|_{1}.$
For each iteration, $\hat{\mathbf{c}}^{k}$ undergoes a gradient descent step to minimize $\mathcal{L}_{\text{text}}$:
$\hat{\mathbf{c}}^{k+1}=\hat{\mathbf{c}}^{k}-\eta_{\text{text}}\cdot\frac{\partial\mathcal{L}_{\text{text}}(\hat{\mathbf{z}}_{t}^{k},\hat{\mathbf{c}}^{k})}{\partial\hat{\mathbf{c}}^{k}},$ (4)
@@ 4.2.3 Point Tracking
After the latent vector $\hat{\mathbf{z}}_{t}^{k}$ and text embedding $\hat{\mathbf{c}}^{k}$ are optimized at $k$-th iteration step, the positions of the handle point $h_{i}^{k}$ should change according to the content dragging. Thus, it is necessary to track the new handle point $h_{i}^{k+1}$ in the updated feature map $\mathcal{F}(\hat{\mathbf{z}}_{t}^{k+1},\hat{\mathbf{c}}^{k+1})$. To find the new handle points $h_{i}^{k+1}$ corresponding to the previous handle points $h_{i}^{k}$, we find the region in the feature map $\mathcal{F}(\hat{\mathbf{z}}_{t}^{k+1},\hat{\mathbf{c}}^{k+1})$ that is most similar to the region around the initial handle points $h_{i}^{0}$ in the feature map $\mathcal{F}(\mathbf{z}_{t}^{0},\hat{\mathbf{c}}^{0})$:
$h_{i}^{k+1}=\mathop{\arg\min}_{q_{2}}\left\|F_{q_{2}}(\hat{\mathbf{z}}_{t}^{k+1},\hat{\mathbf{c}}^{k+1})-F_{h_{i}^{0}}(\mathbf{z}^{0}_{t},\mathbf{c}^{0})\right\|_{1},$ (5)
where $q_{2}$ denotes a point within the region $\Omega(h_{i}^{k},r_{2})$. It ensures that handle points are consistently updated via nearest neighbor search for subsequent optimization iterations.
@@ 5 Experiments
To evaluate the effectiveness of DragText, we conducted a series of experiments on DragDiffusion [31], FreeDrag [13], DragNoise [14], and GoodDrag [42]. The proposed method is applicable to these without hyperparameter tuning, ensuring robustness. Moreover, text optimization hyperparameters are uniformly applied across all methods.
Specifically, we set $\lambda_{\text{text}}=0.1$ to match the image regularization factor $\lambda_{\text{image}}=0.1$, and the text optimization learning rate $\eta_{\text{text}}=0.004$. Detailed information about implementation is in Appendix D. Our method requires no additional GPU resources beyond those used by baseline methods and is executed using a single NVIDIA RTX A6000. Qualitative results are based on DragDiffusion, and further examples using other methods are in Appendix E.
Table 1: Quantitative results on the DragBench dataset [31]
@@
Figure 7: Quantitative results comparing the product of LPIPS and MD for baseline methods and w/ DragText. Performance improved across all methods when applying DragText.
5.1 Qualitative Evaluation
Figure 8: Ablation study on the block number of U-Net decoder, which determines the location where the feature map is obtained. Lower U-Net blocks (e.g., Block 1) focus on low semantic information, while higher U-Net blocks (e.g., Block 4) concentrate on high semantic information. The best results are observed in block 3, where both types of information are appropriately mixed.
In Fig. 5, we compared qualitative results of methods with and without DragText on the DragBench dataset [31]. Existing methods struggled to reach target points or suffer from the degradation of important image semantics, due to the use of image and text embeddings that are not strongly coupled. In contrast, by applying DragText, images can be accurately dragged to the specified target points (e.g., (b) mountain peak and (d) sculpture) while preserving essential semantics that should not disappear (e.g., (c) the mouth of a crocodile). Furthermore, DragText effectively manages content removal and creation caused by dragging, producing high-quality results (e.g. (a) the tongue of a dog and (e) the horns of a goat).
@@ 5.2 Quantitative Evaluation
In Table 1, we present the quantitative comparison of methods with and without DragText on the DragBench dataset [31]. The evaluation metrics are LPIPS [40] and mean distance (MD) [22]. LPIPS measures the perceptual similarity between the edited image to the original image. MD measures the distance between the final position of the handle point $h$ and the target point $g$ in the feature map of the DIFT [35]. While baseline methods and DragText applied methods show little differences in LPIPS, there are significant improvements in MD. As shown in Fig. 7, differences are more dramatic, when considering LPIPS and MD simultaneously. Applying DragText results in substantial performance improvements across all methods, indicating that text embedding optimization allows for effective dragging while preserving image fidelity.
@@ 5.3 Ablation Study
Effect of Text Regularization. In Fig. 6, we investigated the impact of $\lambda_{\text{text}}$ designed to preserve the semantics of the image and text. We edited images by varying $\lambda_{\text{text}}$ from 00 to $10$, which controls the degree of regularization for the text embedding. As $\lambda_{\text{text}}$ approaches 00, the force of dragging becomes more dominant than the preservation of semantics. For example, in Fig. 6 (d), the icecap has disappeared, indicating extensive dragging. Conversely, applying an excessively high $\lambda_{\text{text}}$ during text embedding optimization maintains the original content of the image and inhibits editing. In Fig. 6 (a), minimal dragging is observed, showing that the original image remains largely unchanged. Our experiments showed that setting $\lambda_{\text{text}}$ to $0.1$ achieves the optimal balance between dragging and content preservation. This value can be adjusted based on users’ editing preferences.
Figure 9: Changes in images when manipulating the image and text embeddings after optimization finished. $\omega$ is a factor of the linear interpolation between the original embeddings and the optimized embeddings. As $\omega$ approaches 00, the image resembles the original image; as $\omega$ approaches $1$, the image resembles the dragged image. DragText also generates images outside the original editing range (e.g., extrapolating to $\omega=-1.0$ and $\omega=2$) in the drag direction, preserving the semantic content at the same time.
Effect of the Block Number of U-Net Decoder. In Fig. 8, we perform DragText using feature maps from four different U-Net decoder blocks to assess the impact of each block on text embedding. The image embedding used the feature map from the 3rd block, as in DragDiffusion  [31]. Images optimized using the feature map from the 3rd block exhibited optimal semantic preservation and effective dragging. Feature maps from lower blocks (e.g. Block 1) are hard to maintain semantics, whereas feature maps from higher blocks (e.g. Block 4) preserve semantics well but result in poor dragging performance. This phenomenon is likely due to the lower blocks of the U-Net containing low-frequency information of the image [32]. Quantitative results in Appendix H support our qualitative evaluation.
@@ 6 Manipulating Embeddings
DragText enables the model to comprehend the degree and direction of changes in the image while preserving the important semantic content. In Fig. 9, we applied the same linear interpolation to both the image and text embeddings before and after editing to generate intermediate images. The image and text embeddings of intermediate images are defined as follows: $\hat{\mathbf{z}}=(1-\omega)\mathbf{z}_{0}+\omega\hat{\mathbf{z}}_{t}^{k},\quad\hat{\mathbf{c}}=(1-\omega)\mathbf{c}^{0}+\omega\hat{\mathbf{c}}^{k}.$ When $\omega=0$, the result corresponds to the original image, and as $\omega$ approaches $1$, it gets closer to the dragged image. Remarkably, this phenomenon is maintained even for $\omega\notin[0,1]$, allowing the degree of dragging to be adjusted even beyond the editing process. With DragText, simple manipulations preserved semantic content while allowing control over the degree and direction of the drag, as DragText optimizes the text and image embedding together. In contrast, the baseline model failed to maintain semantics as the text embedding necessary for semantic preservation is not jointly optimized.
@@ 7 Limitation
Diffusion-based editing methods have the common problem of often losing the dragged feature in the latent space, leading to the vanishing of content in surrounding objects that are not the primary editing target. Specifically, in Fig. 10 (a), grape grains within the red circle disappeared when images are dragged, regardless of whether DragText is applied. Nevertheless, this situation changes dramatically when the user provides a more detailed text prompt. For instance, in Fig. 10 (b), when the user provides the exact word “grapes” related to the editing target, DragText dragged the grape feature since it optimizes the text embedding along with the image embedding, preserving the original content. The optimized text embedding plays a complementary role for the features lost in the image embedding space. On the other hand, the baseline model still could not drag the grape feature.